OpenStudy (sreekar369):

if 3x+3y-1,4x2+y-5,4x+2y are the sides of an equilateral triangle ,its area is closet to the integer a)84 b)85 c)86 d)87

OpenStudy (sreekar369):

i want explanasion

OpenStudy (anonymous):

its c i took this test question

OpenStudy (sreekar369):

how it came @DoveOliviaCameron

OpenStudy (sreekar369):

@mayankdevnani

OpenStudy (michele_laino):

@sreekar369 I think you have to provide some numerical data, please

OpenStudy (sreekar369):

they didnt given any other information,they asked to find area@Michele_Laino

OpenStudy (michele_laino):

@sreekar369 are you sure that the equations you provided are corrected?

OpenStudy (sreekar369):

yes

OpenStudy (michele_laino):

@sreekar369 your area is: \[84.77\approx 85\]

OpenStudy (sreekar369):

how u got it @Michele_Laino

OpenStudy (michele_laino):

I't simple, what you have are the measures of the sides of the equlateral triangle. So those measures have to be equal each other, so you have to solve the two equations below: \[3x+3y-1=4x+2y\] \[4x ^{2}+y-5=4x+2y\] if you solve the above system for y, you will get two values for y, but note that only one value of y is acceptable. I will give the acceptable solution: \[y=3, x=2\] other solution for y, namely: \[y=1/4, x=-3/4\] is not acceptable. Now, with the right x and y values, you have to substitute them in your three equations, which are the measures of the sides of your triangle. the mesure of the side of your equilateral triangle is 14. FInally the area of your triangle is: \[14*\frac{ \sqrt{3} }{ 2 }*14*\frac{ 1 }{ 2 }=84.77\]

OpenStudy (sreekar369):

ok thqs

OpenStudy (michele_laino):

thanks!

OpenStudy (sreekar369):

i kept another question once u see it @Michele_Laino

OpenStudy (michele_laino):

Ok!

OpenStudy (michele_laino):

@sreekar369 what is your new question, please?

OpenStudy (sreekar369):

i will tag u in tht